# Image-to-Video Transfer Learning based on Image-Language Foundation Models: A Comprehensive Survey

### v1

@@ VI-H Open Vocabulary Multi-Object Tracking
Results and analysis. In the OVMOT task, researchers explored multiple image-to-video transfer paradigms, including distillation, adapter modules, and post-network tuning, while also experimenting with two distinct foundation models: CLIP[1] and GroundingDINO [4]. As shown in Table III, no single method clearly outperforms the others or achieves superior performance across all metrics. Specifically, the CLIP-based GLATrack model [32], which performs image-to-video transfer via knowledge distillation, achieves the highest scores on TETA and AssocA, surpassing other competitors by more than 1%. In contrast, the GroundingDINO-based approaches MASA [107] and NetTrack [40] perform better on LocA and ClsA, respectively, highlighting that different methods excel in different evaluation criteria.
VII FUTURE DIRECTIONS AND DISCUSSION
@@
This section presents the challenges and potential research directions related to transfer useful knowledge from pre-trained vision-language foundation model for video-text multi-modality understanding.
VII-A Challenges
A key challenge in transferring vision-language foundation models to the video domain lies in the limited and specialized nature of their capabilities. For instance, CLIP is primarily designed for image-text alignment, BLIP focuses on vision-language pre-training for image captioning and question answering, MDETR emphasizes grounding textual descriptions in object detection tasks, while GroundingDINO targets phrase grounding with strong detection performance. Despite their effectiveness within individual domains, no unified model has yet demonstrated robust and comprehensive capabilities across diverse tasks. As a result, different application scenarios in video generation, understanding, or editing often require selecting distinct foundation models, leading to fragmented solutions. Moreover, a standardized and effective strategy for fine-tuning these models on video tasks has yet to be established. Current research predominantly investigates diverse transfer paradigms in an exploratory fashion, highlighting both the immaturity of existing approaches and the experimental stage of this research direction.
VII-B Future Directions
@@
As the transfer of image-language foundation models to video-language tasks continues to evolve, several promising research directions emerge. Below, we outline three key trends that are likely to shape future efforts in this field.
Unified Transfer Learning Paradigm for Multiple Tasks. A significant and promising future direction is the development of a unified transfer learning framework that enables a single ILFM to be effectively adapted for multiple video-language tasks simultaneously. Current methods often require designing distinct fine-tuning strategies or adding task-specific modules for each target task (e.g., action recognition, temporal grounding, video QA), which is inefficient and lacks generalization. Future research should focus on creating a more universal adaptation mechanism. This could involve exploring prompt-based learning, where task-specific instructions guide the model’s behavior, or designing parameter-efficient tuning methods (e.g., unified adapters, hypernetworks) that share most parameters across tasks. The ultimate goal is a single, unified model that can perform a wide array of video understanding tasks through a simple and consistent interface, moving beyond specific solutions toward a more general-purpose video intelligence system.
Collaboration of Multiple Foundation Models. Another promising avenue involves effectively integrating multiple pre-trained foundation models—each specializing in a different modality or capability—to tackle a single video-language task. For example, combining large language models (LLMs) with vision foundation models could enhance contextual reasoning and temporal understanding in complex video tasks. Research efforts could focus on fusion mechanisms, knowledge distillation, and cross-model attention to harness complementary strengths while mitigating computational costs.
Advanced Fusion Methods for Dual-Modality Integration. Despite progress, effectively aligning and fusing visual and linguistic features across spatial and temporal dimensions remains challenging. Future studies may investigate more dynamic and efficient fusion techniques, such as cross-modal transformers, graph-based alignment, or lightweight feature interaction modules. Improving the synergy between video and text modalities will be critical for tasks requiring fine-grained understanding, such as spatio-temporal grounding or long-form video reasoning.
@@
These directions highlight the need for more efficient, integrative, and scalable approaches to bridge image-language models and video-language applications, paving the way for richer video understanding systems.
VII-C More Discussion
This survey primarily reviews the literature on transferring knowledge from image–text foundation models to advance video–text understanding. Indeed, recent studies have increasingly explored extending these foundation models with image-to-video transfer learning techniques for video editing and generation. A brief overview of related works is provided here, and readers are referred to [206] for a more comprehensive and systematic survey.
Methods. Video generation aims to create visually coherent and semantically meaningful video sequences from various inputs, such as text or images. The advent of vision–language foundation models such as Stable Diffusion (SD) [75] has stimulated extensive research on video generation, a task that demands effective modeling of temporal coherence and dynamic motion. Early works, like ModelScope Text-to-Video [85], modified the latent diffusion U-Net by adding 3D convolutions and temporal attention to the frozen spatial attention block, enabling direct text-to-video synthesis in the latent space. At the same time, commercial systems such as Runway Gen-2 [207] began to demonstrate advanced text-to-video capabilities, likely based on optimized spatiotemporal latent diffusion frameworks. A different approach was proposed by AnimateDiff [208], which introduced a lightweight Motion Module trained independently and injected into frozen U-Nets of pre-trained SD checkpoints using Motion LoRA. This design decouples the learning of appearance from that of motion, thereby improving the efficiency of animation. Subsequent works like VideoCrafter1 [83] further refined spatiotemporal U-Net architectures by partial tuning, often using causal 3D convolutions. For image-to-video generation, I2VGen-XL [76] employed a hierarchical latent diffusion framework that upsampled videos progressively across multiple stages. More recently, Stability AI released Stable Video Diffusion (SVD) [74], which extends and fully fine-tunes image-based Stable Diffusion models by incorporating temporal layers. This design enables the generation of short, temporally consistent video clips and even multi-view outputs.
Text-driven video editing aims to generate high-quality videos aligned with a target prompt while preserving the source video’s spatio-temporal structure. The main challenge is balancing flexible semantic edits with temporal consistency. Significant advancements in video editing stem from leveraging large-scale, pretrained vision-language foundation models, particularly Stable Diffusion (SD)[75]. This marks a departure from traditional video processing pipelines that handled frames individually, often leading to flickering artifacts and quality degradation. By adapting models initially trained on static images, researchers now explicitly address the temporal dimension inherent in video data, aiming for coherent edits across sequences.
Pioneering work Tune-A-Video [62] transferred pretrained Stable Diffusion to video tasks through side-tuning spatio-temporal attention modules. It introduced sparse-causal and temporal attention layers within the U-Net to improve consistency, but required costly per-video fine-tuning. To overcome this limitation, Text2Video-Zero [209] proposed a training-free approach. It integrated motion dynamics using cross-frame attention and latent warping, though performance dropped on complex motions due to reliance on implicit cues. Subsequent zero-shot methods explored richer conditioning strategies. Runway Gen-1 [73] combined CLIP embeddings and depth maps with a finetuned convolution–attention U-Net. Pix2Video [134] applied DDIM inversion and depth-guided denoising, preserving structure by injecting latent features from earlier frames without training. ControlNet [136] was adapted in ControlVideo [210], which combined cross-frame attention and interpolation for multimodal structural guidance in a training-free manner. Rerender A Video [211] improved temporal smoothness with sparse causal attention, latent warping, and optical-flow–based fusion in both latent and pixel spaces. More recent works emphasize explicit motion and correspondence modeling. TokenFlow [135] introduced a key-frame–centric global attention strategy. FLATTEN [212] propagated pixel correspondences via optical flow within a training-free temporal-guided attention. Building on this, FRESCO [82] refined flow-based methods with occlusion masks and combined cross-frame flow attention with intra-frame similarity. Together, these developments highlight a shift from implicit motion cues toward explicit modeling, enabling more faithful and temporally coherent video editing.
Evaluation Datasets and Metrics. Video generation and editing rely heavily on large-scale datasets and comprehensive evaluation protocols. Commonly used datasets include WebVid [213] for text–video pairs, LAION [214] for large-scale image–text pairs, Kinetics [6] for human actions, HowTo100M [215] for instructional videos, and DAVIS for video editing benchmarks. These datasets provide rich motion, semantic, and structural priors essential for training and evaluation. System performance is typically assessed using objective metrics such as Fréchet Video Distance (FVD) [216] for temporal consistency and perceptual quality, Fréchet Inception Distance (FID) [217] and Inception Score (IS) [218] for fidelity, and CLIP-based measures [219] for text–video alignment. For editing tasks, frame-wise accuracy (Fram-Acc) and temporal coherence (Tmp-Con) are further employed to evaluate spatial preservation and temporal stability. Beyond these automated measures, human evaluation remains indispensable for judging higher-level aspects such as narrative coherence, creativity, and artifact detection.
Analysis. Recent studies have shown that temporal coherence can be enhanced by integrating temporal attention into diffusion backbones such as Stable Diffusion, either through partial fine-tuning or by incorporating information from preceding frames in a training-free manner. Furthermore, leveraging auxiliary modalities such as depth maps and optical flow strengthens temporal modeling and enables finer-grained spatial understanding, thereby improving the quality of both video generation and editing.
@@ VII-C More Discussion
However, with the growing availability of open-source video foundation models such as SVD [74] and WAN [220], recent studies have shifted toward training and fine-tuning video-based foundation models, which have demonstrated superior performance as compared with image-to-video transfer methods built upon image–language foundation models.
VIII CONCLUSION
@@
This survey presents a comprehensive overview of recent advancements in image-to-video transfer learning within multimodal learning, which, to the best of our knowledge, is the first of its kind. We begin by outlining the necessary background, including fundamental concepts and an introduction to common image-language foundation models (ILFMs). Subsequently, we summarize nine distinct transfer paradigms employed for image-to-video transfer learning. For each video task, we categorize existing ILFM-based methods according to their technical perspectives and the paradigms used for transferring knowledge from ILFMs. In the experimental section, we detail the evaluation settings and present a fair comparison of results to analyze the performance of different transfer paradigms across various tasks. Finally, we discuss several prevailing challenges and suggest promising future research directions for image-to-video transfer learning.

### v2

@@ VII-A Open Vocabulary Multi-Object Tracking
Results and analysis. In the OV-MOT task, researchers explored multiple image-to-video transfer paradigms, including distillation, adapter modules, and post-network tuning, while also experimenting with two distinct foundation models: CLIP[146] and G-DINO [117]. As shown in Table II, no single method clearly outperforms the others or achieves superior performance across all metrics. Specifically, the CLIP-based GLATrack model [86], which performs image-to-video transfer via knowledge distillation, achieves the highest scores on TETA and AssocA, surpassing other competitors by more than 1%. In contrast, the G-DINO-based approaches MASA [98] and NetTrack [212] perform better on LocA and ClsA, respectively, highlighting that different methods excel in different evaluation criteria.
VII-B Spatio-Temporal Video Grounding
Datasets and Metrics. For the STVG task, VidSTG [210] is currently the most widely-used benchmark dataset. It contains 99,943 video-text pairs, including 44,808 declarative sentences and 55,135 interrogative sentences from 10,303 videos spanning 80 object categories. For clearer comparison, we evaluate different methods under the declarative sentence setting. Following [55, 56] we use mean vIoU as the main metric. vIoU is computed as: $vIoU=\frac{1}{\left|T_{u}\right|}\sum_{t\in T_{i}}IoU(\hat{b}_{t},b_{t}),$ where $T_{i}$ and $T_{u}$ denote the intersection and union of the annotated and predicted time intervals, and $\hat{b}_{t},b_{t}$ are the predicted and ground truth bounding boxes for the $t$-th frame, respectively. We also report vIoU@R which indicates the proportion of samples with vIoU higher than $R$. Following [195], we report tIoU evaluating the temporal grounding performances. It is defined as: $tIoU=\frac{\left|T_{i}\right|}{\left|T_{u}\right|}.$
Results and analysis. The comparative results on VidSTG-Declarative[210] are shown in Table II. Most fully-supervised STVG methods adopt MDETR as the foundation model for its visual grounding capability. However, relying solely on its image-level features limits temporal localization accuracy. Recent methods therefore integrate auxiliary models (e.g., SlowFast[40]) to extract video-level features, improving temporal performance—tIoU rose from 48.1 (TubeDETR[195]) to 52.1 (CoSTA[105]). This indicates that combining both feature types enables more precise spatio-temporal localization, though at increased computational cost (e.g., TASTVG[56] required 32 A100 GPUs for 15 hours).
VII-C Temporal Video Grounding
Datasets and Metrics. To address the TVG problem, numerous datasets have been constructed with video–text pairs and detailed annotations specifying the temporal spans of events described in the text. Specifically, TACoS [148], Charades-STA [44] and ActivityNet Captions [79] are the three most widely used benchmarks in the community. Following [44], most existing approaches are evaluated in Charades-STA [44] and report their performance in terms of metrics “Rn@m” and “mIoU”. Metric “Rn@m” is employed to compute the percentage of samples where the top-$n$ predicted segments have an IoU greater than a predefined threshold $m$, where $m$ is typically defined as $\{0.1,0.3,0.5,0.7\}$. Metric “mIoU” calculates the mean IoU between the predicted segments and ground truth on all testing data.
Results and Analysis. We report comparison results on the Charades-STA [44] set with “Rn@m” and “mIoU” metrics. As shown in Table II, the results for the TVG task reveal a clear performance gap between different foundation models. Previously, the CLIP-based method R2-tuning [118] achieved the best performance with 50.9% on mIoU, surpassing LLaVA-based methods like NumPro[185] with 41.4% mIoU. It is worth noting that NumPro relies on a complex Fine-Tuning with Extra Models strategy, while R2-tuning uses Side-Tuning based on MLLM. However, the recent VideoTG-R1[33] built on Qwen2.5-VL obtains the highest scores across all metrics, reaching 55.9% in the term of mIoU. Surprisingly, VideoTG-R1 employs a simple Post Network Tuning strategy but outperforms all previous methods. This demonstrates that a stronger foundation model can achieve superior temporal localization without complex auxiliary modules.
VII-D Video-Text Retrieval
Datasets and Metrics. To facilitate VTR research, several datasets pairing videos with text descriptions have been created to evaluate cross-modal alignment. Among these, MSR-VTT [200] is a widely used benchmark. Following common practice, most methods are evaluated on this dataset and report R@1, R@5, R@10, and R@sum. R@1, R@5, and R@10 measure recall among the top 1, 5, and 10 results, reflecting ranking accuracy. R@sum aggregates these three scores to summarize overall retrieval performance.
Results and analysis. By examining the results on MSR-VTT [200] in Table III for the task of Video-Text Retrieval, it can be observed that most recent methods are built upon the pretrained CLIP model due to its strong visual-textual correspondence capability. Among these, fine-tuning with LoRA achieves the best performance, significantly surpassing other feature-modification transfer learning strategies such as prompt tuning, knowledge distillation, and adapter-based tuning, in the metric of R@1. As expected, directly fusing CLIP-extracted features with a post-network tuning scheme performs worse than most approaches that update CLIP parameters, highlighting the importance of adapting the original foundation model to the target task and dataset for optimal performance.
VII-E Video Action Recognition
Datasets and Metrics. To support research in VAR, several influential datasets have been developed. Foundational benchmarks include UCF101 [154] and HMDB51 [80], which established early standards for the task. More recently, the large-scale Kinetics series Kinetics-400 [76] and Something-Something V2 [54] have become the most widely adopted benchmarks for evaluating VAR models. Following common practice, model performance on these datasets is primarily evaluated using Top-1 and Top-5 accuracy. Top-1 accuracy measures the percentage of videos for which the model’s single highest-probability prediction is correct. Top-5 accuracy measures the percentage of videos where the ground-truth label is among the model’s five most confident predictions, which is particularly informative for datasets with a large number of classes.
Results and analysis. Table III (termed as VAR) reports a detailed comparison of several representative image-to-video transfer methods for video action recognition based on the CLIP model. Overall, finetuning with Adapter strategy achieves better results than side-tuning strategy in this task. MoMa [198] obtains the best results on both K400 and SSv2 sets, which employs a mamba architecture to learn long-term fine-grained spatio-temporal information. The comparison between EVL [111] and DiST [145] demonstrates the effectiveness of disentangling the spatial and temporal modeling in image-to-video transfer learning.
VII-F Video Question Answering
Datasets and Metrics. VideoQA is typically evaluated using metrics such as Top-1 Accuracy and Top-K Accuracy. The VideoQA task typically comprises three categories: i) Multiple-Choice: Requires selecting the correct answer from a predefined set. Representative benchmarks such as MSRVTT-MC [200]. ii) Open-Ended: Involves generating free-form responses that often necessitate logical or creative reasoning. Popular datasets include MSRVTT-QA [200], MSVD-QA [188], and ActivityNet-QA [201]. iii) Long-Form: Demands comprehensive, explanatory answers reflecting deep temporal understanding, with key benchmarks such as How2QA [94].
Results and analysis. We present the comparison results on four datasets (MSRVTT-MC[200], NExT-QA[186], MSVD-QA[188] and How2QA[94]) with Top-1 Accuracy on Multiple-Choice in Table III, where most competitors use CLIP as their backbone to ensure a fair comparison. For standard vision-language models such as CLIP, various transfer methods demonstrate strong and dataset-dependent performance. The distillation-based approach, VideoDistill [219], achieves SOTA results on MSRVTT-MC and MSVD-QA with accuracies of 97.8% and 49.2%, respectively. In contrast, on the How2QA benchmark, the prompt-tuning method Q-Vid [133] attains the highest score at 71.4%. These results suggest that for a fixed backbone like CLIP, the most effective adaptation strategy varies across different benchmarks. Conversely, the integration of LLMs substantially elevates performance. On the NExT-QA benchmark, the LLaVA-based BIMBA [68] model, which utilizes a post-network tuning strategy, achieves an accuracy of 83.73%. This represents a significant improvement over the best CLIP-based method, Q-Vid with an accuracy of 66.3% on the same dataset. This finding underscores that post-network tuning is a potent transfer strategy for LLM-based models, as it efficiently adapts visual information to unlock the LLM’s powerful semantic reasoning for video understanding.
VII-G Video Captioning
Datasets and Metrics. Prominent open-domain datasets like MSRVTT [191] and MSVD [15] assess general descriptive capabilities, while procedural datasets such as YouCook2 [214] challenge models with fine-grained, instructional content. Performance on these benchmarks is typically quantified by comparing generated captions against human references using a suite of automated metrics. Specifically, BLEU@N [139] and ROUGE-L [107] focus on lexical similarity and recall, whereas METEOR [6] and CIDEr [166] provide a more semantically informed evaluation by considering synonyms and consensus, collectively ensuring a holistic assessment of linguistic fluency and content fidelity.
Results and analysis. As shown in Table III, recent advancements in video captioning are driven by adapted transfer strategies that outperform traditional fine-tuning. The side-tuning method MAMS [84] achieves SOTA performance across multiple key metrics (60.0 BLEU4, 71.2 ROUGE-L, 82.9 CIDEr). This success can be primarily attributed to its architectural design; by training a separate, lightweight module while keeping the foundation model frozen, it effectively learns task-specific temporal dynamics without the risk of catastrophic forgetting. The strong performance of the LLaVA-based HierarQ [2], which excels by aligning visual features with an LLM’s powerful reasoning, further validates this trend. These results indicate that the most effective transfer strategies are those that intelligently integrate new knowledge while preserving the core capabilities of the pre-trained backbone.
VIII FUTURE DIRECTIONS AND DISCUSSION
@@
In this section, we aim to give some insights on key challenges and potential future research directions related to transferring useful knowledge from pre-trained image-language foundation models for video-text multimodal understanding.
VIII-A Challenges and Future Directions
A principal challenge in adapting vision-language foundation models to video understanding stems from their inherently specialized architectures. For instance, CLIP is designed for image-text alignment, while MDETR and G-DINO focus on visually-grounded object detection. More recent models, such as LLaVA, InternVL, and QwenVL, have extended these capabilities toward general-purpose visual reasoning and dialogue, yet they remain fundamentally designed for the image domain.
Despite their individual strengths, no single model has yet served as a unified solution that performs robustly and comprehensively across diverse video tasks. As a result, video understanding systems often rely on fragmented pipelines that combine separate foundation models for different sub-tasks. Moreover, how to effectively fine‑tune image‑based models—from established ones like CLIP to more modern multimodal LLMs such as LLaVA and InternVL—on sequential video data remains an open research question.
@@ VIII-A Challenges and Future Directions
As the transfer of ILFMs to video-language tasks continues to evolve, several promising directions emerge. Below, we outline three key trends that are likely to shape future efforts.
Unified Transfer Learning Paradigm for General Video Understanding. Since existing ILFMs excel at specific tasks, their cross-modal capabilities can be transferred to address certain video-language understanding problem, such as MDETR for video grounding. A promising future direction involves developing a unified transfer learning framework to transfer a single ILFM for multiple general video-language understanding tasks simultaneously. This can be achieved by training a unified ILFM capable of performing general image-language understanding, or by prioritizing the design of more universal transfer mechanisms. Potential approaches include exploring prompt-based learning paradigms, where task-specific instructions guide model behavior; incorporating video-oriented reinforcement learning to enhance comprehension; and developing parameter-efficient tuning strategies that facilitate parameter sharing across diverse tasks.
Collaborative Transfer of Multiple Foundation Models. Beyond the development of a unified ILFM, another promising direction lies in the effective integration of multiple pre-trained foundation models, each excelling in a distinct modality or capability, to address complex video-language tasks. Within this paradigm, non-LLM-based models can function as external tools within a video understanding pipeline, generating intermediate outputs such as per-frame spatial bounding boxes from models like MDETR or G-DINO. Concurrently, LLM-based models can serve as central agents, performing reasoning-based inference and orchestrating these tools. Future research should focus on fully leveraging the potential of LLMs in this agentic role while developing methodologies to mitigate the associated computational overhead.
Advanced Efficient Methods for Video Frame Processing. While significant progress has been made, efficiently leveraging ILFMs for video frame processing remains an open challenge. Existing adaptation approaches predominantly rely on uniform frame sampling strategies, which are suboptimal due to the inherent high degree of visual redundancy in video sequences. Consequently, a critical research direction involves designing adaptive frame sampling mechanisms that minimize the number of frames processed by ILFMs without compromising task-critical information. Furthermore, enhancing local alignment between video and text modalities is essential for advancing tasks that demand fine-grained understanding, such as spatiotemporal grounding or long-form video reasoning.
@@
These directions highlight the need for more efficient, integrative, and scalable approaches to bridge image-language models and video-language applications, paving the way for more efficient and diverse video understanding systems.
VIII-B More Discussion
These future directions mainly target video–language understanding, aiming to make ILFMs more general, collaborative, and efficient under video input streams. Beyond understanding, a parallel and rapidly growing line of research investigates transferring generative ILFMs, such as Stable Diffusion (SD) [151], to video generation and editing. For example, VideoCrafter1 [17] further refined spatiotemporal U-Net architectures by partial tuning, often using causal 3D convolutions. Stability AI released Stable Video Diffusion (SVD) [8], which extends and fully fine-tunes image-based Stable Diffusion models by incorporating temporal layers. Tune-A-Video [181] transferred pretrained SD to video tasks through side-tuning spatio-temporal attention modules.
@@ VIII-B More Discussion
More recent studies have shifted toward training and fine-tuning video-based foundation models, such as SVD [8] and WAN [167], which have demonstrated superior performance as compared with image-to-video transfer methods built upon image–language foundation models. Readers can refer to  [61, 156, 132] for a more comprehensive and systematic survey.
IX CONCLUSION
@@
This survey presents a comprehensive overview of recent advancements in image-to-video transfer learning within multimodal learning, which, to the best of our knowledge, is the first of its kind. We begin by outlining the necessary background, including fundamental concepts and an introduction to common ILFMs. Subsequently, we summarize nine distinct transfer paradigms employed for image-to-video transfer learning. For each video task, we categorize existing ILFM-based methods according to their technical perspectives and the paradigms used for transferring knowledge from ILFMs. In the experimental section, we detail the evaluation settings and present a fair comparison of results to analyze the performance of different transfer paradigms across various tasks. Finally, we discuss several prevailing challenges and suggest promising future research directions for image-to-video transfer learning.